OpenStudy (anonymous):

The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46 centimeters. The measure y is 4 centimeters longer than the measure x. Write two equations which represent the two sentences. and the second part of the question The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x. Write two equations which represent the two sentences. Solve the system you wrote for this using any technique.

OpenStudy (anonymous):

y=x+4 x+x+y=46

OpenStudy (anonymous):

can you do this part? Solve the system you wrote for this using any technique.

OpenStudy (anonymous):

put relation for y from above in next equation 3x+4=46 3x=42 x=14

OpenStudy (anonymous):

thanks